# 3D Beamforming Through Joint Phase-Time Arrays

### v1

@@ III-A2 Infinity Norm Solution
In this part, we define the problem as $\min_{\bm{\tau},\bm{\phi}}\lVert\mathbf{e}_{i}(\bm{\theta}_{\rm az},\bm{\theta}_{\rm el},\mathcal{F}_{m})\rVert_{\infty}$. However, this approach does not lead to a closed-form solution and the optimization can be solved by using linear programming toolbox, for example, CVX.
@@ III-B Separated Analytical Solution
We can rewrite the antenna array steering vector, which is given in Eq. (1), by the Kronecker product of azimuth and elevation antenna array steering vectors as in
$\mathbf{a}_{\rm R}(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=\mathbf{a}_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})\otimes\mathbf{a}_{\rm el}(\theta_{\rm el},f_{m}),$ (14)
$\mathbf{a}_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=e^{-j\pi\frac{f_{m}}{f_{c}}y\sin\theta_{\rm az}\sin\theta_{\rm el}},$
$\mathbf{a}_{\rm el}(\theta_{\rm el},f_{m})$ $=e^{-j\pi\frac{f_{m}}{f_{c}}z\cos\theta_{\rm el}}.$
@@
$\bm{\omega}(y,z,f_{m})$ $=\bm{\omega}_{\rm az}(y,f_{m})\otimes\bm{\omega}_{\rm el}(z,f_{m}),$ (15)
$\bm{\omega}_{\rm az}(y,f_{m})$ $=e^{j({\phi_{{\rm az},y}}+2\pi f_{m}{\tau_{{\rm az},y}})},$
$\bm{\omega}_{\rm el}(z,f_{m})$ $=e^{j({\phi_{{\rm el},z}}+2\pi f_{m}{\tau_{{\rm el},z}})}.$
By using the above descriptions, the beamforming gain is $G(\theta_{\rm az},\theta_{\rm el},f_{m})=G_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})G_{\rm el}(\theta_{\rm el},f_{m})$ for
$G_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=\frac{1}{N_{\rm az}}\left\lVert\sum_{y=0}^{N_{\rm az}-1}e^{jh_{{\rm az},y}-j\Omega_{{\rm az},y}}\right\rVert^{2},$ (16)
with the definitions of $h_{{\rm az},y}\triangleq{\phi_{{\rm az},y}}+2\pi f_{m}{\tau_{{\rm az},y}}$ and $\Omega_{{\rm az},y}\triangleq\pi\frac{f_{m}}{f_{c}}y\sin\theta_{\rm az}\sin\theta_{\rm el}$, and
$G_{\rm el}(\theta_{\rm el},f_{m})$ $=\frac{1}{N_{\rm el}}\left\lVert\sum_{z=0}^{N_{\rm el}-1}e^{jh_{{\rm el},z}-j\Omega_{{\rm el},z}}\right\rVert^{2},$ (17)
with the definitions of $h_{{\rm el},z}\triangleq{\phi_{{\rm el},z}}+2\pi f_{m}{\tau_{{\rm el},z}}$ and $\Omega_{{\rm el},z}\triangleq\pi\frac{f_{m}}{f_{c}}z\cos\theta_{\rm el}$.
The linear system of equations is formed in a similar way to Sec. III-A. However, we individually engage in the solution of $y$-th antenna with the specific objective of optimizing $G_{\rm az}$, and $z$-th antenna with the intent of maximizing $G_{\rm el}$ since these two gain values are inherently independent of one another, as per their definitions. Consequently, this alternate approach requires $N_{\rm az}+N_{\rm el}$ equations while joint configuration necessitates $N_{\rm az}\times N_{\rm el}$ due to individually interacting with each antenna element.
@@ III-C Greedy Algorithm
The objective of the greedy algorithm is to explore the solution space to identify any potential enhancements in the defined gain, Eq. (3), after analytical derivations. Given that we have established two distinct solution sets for $\bm{\tau}$ and $\bm{\phi}$ through the joint and separated optimization techniques, we will subsequently pursue two different greedy optimization approaches: the joint and separated. For both cases, parameters are initialized by using the least squares method because this method results in a closed form solution, therefore; it’s one-shot, whereas the infinity norm requires linear programming so the solution is iterative.
@@ III-C1 Joint
Within the two-dimensional antenna array, our algorithm explores the delay grid for a specified antenna element to improve $G_{l}$ while holding the delay and phase values of other antenna elements constant. Subsequently, a similar process is applied to the phase values. These updates continue until a predefined convergence criterion is satisfied. The criterion is the absolute difference in $G_{l}$ between consecutive runs being smaller than $\zeta$ of the last iteration. We explore the various $\zeta$ values to find the optimal one. Once the convergence criteria are met, the algorithm returns the updated $\bm{\phi}$ and $\bm{\tau}$.
1 Input: $N_{\rm az}$, $N_{\rm el}$, $M$, $\bm{\theta}_{\rm az}$, $\bm{\theta}_{\rm el}$, $\bm{\alpha}$ Output: $\widetilde{\bm{\tau}}$, $\widetilde{\bm{\phi}}$ Initialize $\bm{\phi}$ and $\bm{\tau}$ by using Eqs. (11) and (12) Define the candidate delays as $\tau_{\rm grid}=0:\tau_{p}:\tau_{\max}$ Define the possible phase values as $\phi_{\rm grid}=0:\frac{2\pi}{2^{\beta}}:2\pi$ $\widetilde{\tau}_{y,z}=\operatorname*{arg\,min}_{\tau_{grid}}\lVert\tau_{y,z}-\tau_{\rm grid}\rVert$ $\forall(y,z)$ $\widetilde{\phi}_{y,z}=\operatorname*{arg\,min}_{\phi_{grid}}\lVert\phi_{y_{z}}-\phi_{\rm grid}\rVert$ $\forall(y,z)$ not-conv $\leftarrow$ True while  not-conv  do
2       $G_{l,{\rm first}}$leftarrow Eq. (4) by $\widetilde{\bm{\tau}}$ and $\widetilde{\bm{\phi}}$ for y = 1:1:$N_{\rm az}$ do
@@
9
10      $G_{l,{\rm later}}$leftarrow Eq. (4) by $\widetilde{\bm{\tau}}$ and $\widetilde{\bm{\phi}}$ if $|G_{l,{\rm later}}-G_{l,{\rm first}}|<\zeta\times G_{l,{\rm later}}$ then not-conv=False
@@ III-C2 Separated
Similarly to the approach described above, we update the parameters by searching within the defined space for phase and delay values. However, in this case, we initialize the parameters as $\bm{\tau}_{\rm az}$, $\bm{\phi}_{\rm az}$, $\bm{\tau}_{\rm el}$ and $\bm{\phi}_{\rm el}$, according to Sec. III-B. The update process now handles $N_{\rm az}+N_{\rm el}$ delay and phase values, compared to the prior approach’s $N_{\rm az}\times N_{\rm el}$ updates.
@@ III-D Gradient Descent Algorithm
The greedy algorithm, although guaranteed to converge to a local optimal solution, takes a long time to converge even for the separated case. The primary objective of the gradient descent algorithm is to quickly explore the solution space with the aim of achieving an improved gain, as defined in (3).
We approach the solution by using an iterative optimization technique that updates the parameters according to their gradients from the defined loss function as in
$F(\bm{\tau},\bm{\phi})=\lVert G_{l,\rm max}-G_{l}(\bm{\tau},\bm{\phi})\rVert^{2},$ (18)
where $G_{l\rm,max}\triangleq 10\log_{10}\left(N_{\rm az}N_{\rm el}\right)$ is the maximum beamforming gain. During this optimization, we schedule learning using Adam optimizer with a learning rate of $0.1$, determined through an exploration of various learning rates using grid search. We apply this method to both solution spaces: joint and separated optimizations.
@@ III-D1 Joint
For the joint case, the optimization parameters are $\bm{\phi}$ and $\bm{\tau}$, initialized according to (11) and (12) similar to greedy optimization. After initialization, we quantize these values with a specific precision for delay, $\tau_{p}$, and bit size for phase, $\beta$. Then, these parameters, $\bm{\phi}$ and $\bm{\tau}$, are updated by using their gradients until the convergence criteria are met, which is the same criteria as in Line 18 of Algorithm 1.
@@ IV Simulation Results
Figure 2: Performance comparison of joint and separated analytical derivations by using LS or infinity norm definition with a) various bandwidth allocation scenarios and b) an unequal bandwidth allocation for $N_{u}=5$ and $\bm{\alpha}=[0.3,0.2,0.15,0.1,0.25]$.
@@
Figure 3: Illustration of the maximum beamforming gain in azimuth and elevation domain for a) separated LS, and b) joint LS when $N_{u}=5$ with equal bandwidth allocation.
Refer to caption (a) Separated LS Refer to caption (b) Joint LS
Figure 3: Illustration of the maximum beamforming gain in azimuth and elevation domain for a) separated LS, and b) joint LS when $N_{u}=5$ with equal bandwidth allocation.
@@
Figure 4: Beamforming gain of joint LS across an azimuth vs frequency domains when $\theta_{\rm el}=105^{\circ}$.
@@
Figure 5: Performance comparison of greedy and gradient descent algorithms using joint and separated optimization with a) an equal bandwidth allocation for $N_{u}=4$ users, and b) various bandwidth allocation scenarios for $N_{u}=2$ users.
Refer to caption (a) 4-user, equal bandwidth allocation Refer to caption (b) 2-user, unequal bandwidth allocation
Figure 5: Performance comparison of greedy and gradient descent algorithms using joint and separated optimization with a) an equal bandwidth allocation for $N_{u}=4$ users, and b) various bandwidth allocation scenarios for $N_{u}=2$ users.
@@
Figure 6: Comparison of our proposed solutions with the state-of-the-art over various bandwidth allocation scenarios
@@
In this section, we provide a numerical evaluation of the methods described in Sec. III. The simulation parameters, which are consistent with a practical 5G mmWave deployment, are given in Table I. For simplicity, we assign $i$-th user’s AoA as $(\theta_{{\rm az},i},\theta_{{\rm el},i})=((-60+i\frac{120}{N_{u}-1})^{\circ},(90+i\frac{30}{N_{u}-1})^{\circ})$ when there are $N_{u}$ users in a cell with $120^{\circ}$ horizontal coverage and $30^{\circ}$ vertical coverage. Note that we subject the output of the functions, phase, and delay values, to quantization before illustrating their results. We present our findings of both analytical (Sec. IV-A) and iterative (Sec. IV-B) results.
IV-A Analytical Results
@@
In Secs. III-A and III-B, we present our analytical methods to configure PSs and TTD units. In Fig. 1(a), we illustrate the average gains of the users, which is denoted as $G_{l}$ in (4), for $N_{u}=2$, where User 1 gets allocated $\alpha$ of the total bandwidth (${\rm BW}$), and User 2 gets $1-\alpha$ of the bandwidth. Our findings reveal that, primarily, joint optimization consistently outperforms separated optimization by a margin ranging from $0.6$ dB to $4$ dB, contingent on the bandwidth allocation ratio. The mean difference between joint and separated optimization, observed across various bandwidth allocations, is approximately $1.29$ dB when the method is least squares (LS) and $1.18$ dB when the method is infinity norm. Hence, it can be deduced that the greater number of systems of equations in the analytical derivation proves to be advantageous.
Moreover, we observe that the closed-form solution obtained through LS optimization results in a lower average gain for unequal bandwidth allocation ($\alpha\leq 0.2$) in both joint and separated optimization scenarios. The difference between the average gain is at most $1.81$ dB and on average $0.76$ dB over different bandwidth allocations for the joint and at most $3.36$ dB and on average $1.4$ dB for the separated optimization.
@@
In Fig. 1(b), we demonstrate $G_{{\rm mean},i}$ for $i$th user, defined in Eq. 5, when $N_{u}=5$. The bandwidth allocations of the users are $\bm{\alpha}=[0.3,0.2,0.15,0.1,0.25]$. In joint optimization, the infinity norm method yields a maximum gain difference of $3$ dB among users, whereas the LS method results in a $5$ dB difference. A similar trend is observed in separated optimization, with the infinity norm leading to a maximum difference of $3.2$ dB, while LS results in a $5.4$ dB difference. This suggests that the infinity norm method fosters greater fairness in user gain distribution, even when their assigned ratios are unequal. Furthermore, due to the more equitable distribution of average gains among users, the infinity norm also elevates the average gain of users, as illustrated in Fig. 1(a).
@@
Furthermore, we illustrate the impact of joint and separated optimization on beamforming gain in Fig. 4. Here, we demonstrate the maximum gain among subcarriers for every azimuth and elevation angle when the desired user locations are marked with red diamonds. It is evident that there are more sidelobes observed in the case of separated optimization, as depicted in Fig. 2(a), in contrast to the joint optimization shown in Fig. 2(b). This difference can be attributed to the interdependence of the phase and delay elements. In Fig. 4, we display the gain across various subcarriers while maintaining a constant elevation angle, which corresponds to the User 3 in Fig. 2(b) with $(\theta_{\rm az},\theta_{\rm el})=(0^{\circ},105^{\circ})$. Here, the red line is the desired maximum gain allocation, which is fairly satisfied by using joint optimization.
IV-B Iterative Results

### v3

@@ III-A2 Infinity Norm Solution
In this part, we define the problem as $\min_{\bm{\tau},\bm{\phi}}\lVert\bm{e}_{i}(\bm{\theta}_{\rm az},\bm{\theta}_{\rm el},\mathcal{F}_{m})\rVert_{\infty}$. However, this approach does not lead to a closed-form solution, and the optimization can be solved by using a linear programming toolbox, for example, CVX.
@@ III-B Separated Analytical Solution
We can rewrite the antenna array steering vector, which is given in Eq. (1), by the Kronecker product, $\otimes$ of azimuth and elevation antenna array steering vectors as in
$\bm{a}_{\rm R}(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=\bm{a}_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})\otimes\bm{a}_{\rm el}(\theta_{\rm el},f_{m}),$ (14)
$\left[\bm{a}_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})\right]_{y}$ $=e^{j\pi\frac{f_{m}}{f_{c}}y\sin\theta_{\rm az}\sin\theta_{\rm el}},$
$\left[\bm{a}_{\rm el}(\theta_{\rm el},f_{m})\right]_{z}$ $=e^{j\pi\frac{f_{m}}{f_{c}}z\cos\theta_{\rm el}}.$
@@
$\bm{\omega}(f_{m})$ $=\bm{\omega}_{\rm az}(f_{m})\otimes\bm{\omega}_{\rm el}(f_{m}),$ (15)
$\left[\bm{\omega}_{\rm az}(f_{m})\right]_{y}$ $=\frac{1}{\sqrt{N_{\rm az}}}e^{j({\phi_{{\rm az},y}}+2\pi f_{m}{\tau_{{\rm az},y}})},$
$\left[\bm{\omega}_{\rm el}(f_{m})\right]_{z}$ $=\frac{1}{\sqrt{N_{\rm el}}}e^{j({\phi_{{\rm el},z}}+2\pi f_{m}{\tau_{{\rm el},z}})}.$
By using the above descriptions, the beamforming gain can be expressed as $G(\theta_{\rm az},\theta_{\rm el},f_{m})=G_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})G_{\rm el}(\theta_{\rm el},f_{m})$, where
$G_{\rm az}(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=\frac{1}{N_{\rm az}}\left\lVert\sum_{y=0}^{N_{\rm az}-1}e^{jh_{{\rm az}}(y)-j\Omega_{{\rm az}}(y)}\right\rVert^{2},$ (16)
with the definitions of $h_{{\rm az}}(y)\triangleq{\phi_{{\rm az},y}}+2\pi f_{m}{\tau_{{\rm az},y}}$ and $\Omega_{{\rm az}}(y)\triangleq\pi\frac{f_{m}}{f_{c}}y\sin\theta_{\rm az}\sin\theta_{\rm el}$, and
$G_{\rm el}(\theta_{\rm el},f_{m})$ $=\frac{1}{N_{\rm el}}\left\lVert\sum_{z=0}^{N_{\rm el}-1}e^{jh_{{\rm el}}(z)-j\Omega_{{\rm el}}(z)}\right\rVert^{2},$ (17)
with the definitions of $h_{{\rm el}}(z)\triangleq{\phi_{{\rm el},z}}+2\pi f_{m}{\tau_{{\rm el},z}}$ and $\Omega_{{\rm el}}(z)\triangleq\pi\frac{f_{m}}{f_{c}}z\cos\theta_{\rm el}$.
The linear system of equations is formed similarly to Sec. III-A. However, we individually engage in the solution of $y$-th antenna with the specific objective of optimizing $G_{\rm az}$, and $z$-th antenna with the intent of maximizing $G_{\rm el}$ since these two gain values are inherently independent of one another, as per their definitions. Consequently, this alternate approach requires $N_{\rm az}+N_{\rm el}$ equations. At the same time, joint configuration necessitates $N_{\rm az}\times N_{\rm el}$ due to individually interacting with each antenna element.
@@ III-C Greedy Algorithm
The objective of the greedy algorithm is to explore the solution space to identify any potential enhancements in the beamforming gain, after analytical derivations. Given that we have established two distinct solution sets for $\bm{\tau}$ and $\bm{\phi}$ through the joint and separated optimization techniques, we will subsequently pursue two different greedy optimization approaches: the joint and separated. For both cases, parameters are initialized by using the least squares method because this method results in a closed-form solution.
@@ III-C1 Joint
Within the two-dimensional antenna array, our algorithm explores the delay grid for a specified antenna element to improve $G_{l}$ while holding other antenna elements’ constant delay and phase values. Subsequently, a similar process is applied to the phase values. These updates continue until a predefined convergence criterion is satisfied. The criterion is the absolute difference in $G_{l}$ between consecutive runs being smaller than $\zeta$ of the last iteration. We explored the various $\zeta$ values to find the optimal one. Once the convergence criteria are met, the algorithm returns the updated $\bm{\phi}$ and $\bm{\tau}$.
1 Input: $N_{\rm az}$, $N_{\rm el}$, $M$, $\bm{\theta}_{\rm az}$, $\bm{\theta}_{\rm el}$, $\bm{\alpha}$
2 Output: $\widetilde{\bm{\tau}}$, $\widetilde{\bm{\phi}}$
3 Initialize $\bm{\phi}$ and $\bm{\tau}$ by using Eqs. (11) and (12)
4
5Define the candidate delays as $\tau_{\rm grid}=0:\tau_{p}:\tau_{\max}$
6
7Define the possible phase values as $\phi_{\rm grid}=0:\frac{2\pi}{2^{\beta}}:2\pi$
8
[…]
10 $\widetilde{\phi}_{y,z}=\operatorname*{arg\,min}_{\phi_{grid}}\lVert\phi_{y_{z}}-\phi_{\rm grid}\rVert$ $\forall(y,z)$
11
12not-conv $\leftarrow$ True
13
14while  not-conv  do
15       $G_{l,{\rm first}}$leftarrow Eq. (5) by $\widetilde{\bm{\tau}}$ and $\widetilde{\bm{\phi}}$
16       for y = 1:1:$N_{\rm az}$ do
@@ 2 Output: $\widetilde{\bm{\tau}}$, $\widetilde{\bm{\phi}}$
24
25      $G_{l,{\rm later}}$leftarrow Eq. (5) by $\widetilde{\bm{\tau}}$ and $\widetilde{\bm{\phi}}$
26       if $|G_{l,{\rm later}}-G_{l,{\rm first}}|<\zeta\times G_{l,{\rm later}}$ then not-conv=False
27
@@ III-C2 Separated
Similarly to the abovementioned approach, we update the parameters by searching within the defined space for phase and delay values. However, in this case, we initialize the parameters as $\bm{\tau}_{\rm az}$, $\bm{\phi}_{\rm az}$, $\bm{\tau}_{\rm el}$ and $\bm{\phi}_{\rm el}$, according to Sec. III-B. The update process now handles $N_{\rm az}+N_{\rm el}$ updates, compared to the prior approach’s $N_{\rm az}\times N_{\rm el}$.
@@ III-D Gradient Descent Algorithm
Although the greedy algorithm is guaranteed to converge to a local optimal solution, it takes a long time to do so even for the separated case. The primary objective of the gradient descent algorithm is to quickly explore the solution space to achieve an improved beamforming gain.
We approach the solution by using an iterative optimization technique that updates the parameters according to their gradients from the defined loss function, as in
$F(\bm{\tau},\bm{\phi})\triangleq\lVert G_{l,\rm max}-G_{l}(\bm{\tau},\bm{\phi})\rVert^{2},$ (18)
where $G_{l\rm,max}\triangleq 10\log_{10}\left(N_{\rm az}N_{\rm el}\right)$ is the maximum beamforming gain. During this optimization, we schedule learning using Adam optimizer with a learning rate of $0.1$, which is determined by exploring various learning rates using a grid search. We apply this method to both solution spaces: joint and separated optimizations.
@@ III-D1 Joint
For the joint case, the optimization parameters are $\bm{\phi}$ and $\bm{\tau}$, initialized according to (11) and (12) similar to greedy optimization. After initialization, we quantize these values with a specific precision for delay, $\tau_{p}$, and bit size for phase, $\beta$. Then, these parameters, $\bm{\phi}$, and $\bm{\tau}$, are updated using their gradients until meeting the convergence criteria, which is the same as Line 18 of Algorithm 1.
@@ IV Simulation Results
Figure 2: Performance comparison of joint and separated analytical derivations by using LS or infinity norm definition with a) various BW allocation scenarios and b) an unequal BW allocation for $N_{u}=5$ and $\bm{\alpha}=[0.3,0.2,0.15,0.1,0.25]$.
(a) Separated LS [10]
(b) Joint LS [10]
@@
Figure 3: Illustration of the maximum beamforming gain in azimuth and elevation domain for a) separated LS, and b) joint LS when $N_{u}=5$ with equal BW allocation.
@@
Figure 4: Beamforming gain of joint LS across an azimuth vs frequency domains when $\theta_{\rm el}=105^{\circ}$. User 3 is located at $(\theta_{\rm az},\theta_{\rm el})=(0^{\circ},105^{\circ})$, and assigned the frequency band $[-9.5,9.5)$ MHz.
(a) 4-user, equal BW allocation
(b) 2-user, unequal BW allocation
@@
Figure 5: Performance comparison of greedy and gradient descent algorithms using joint and separated optimization with a) an equal BW allocation for $N_{u}=4$ users, and b) various BW allocation scenarios for $N_{u}=2$ users. Greedy and gradient descent algorithms yield similar performance.
@@
Figure 6: Comparison of our proposed solutions with the state-of-the-art over various BW allocation scenarios
@@
This section provides a numerical evaluation of the methods described in Sec. III. The simulation parameters, which are consistent with a practical 5G mmWave deployment, are given in Table I. For simplicity, we assign $i$-th user’s AoA as $(\theta_{{\rm az},i},\theta_{{\rm el},i})=((-60+i\frac{120}{N_{u}-1})^{\circ},(90+i\frac{30}{N_{u}-1})^{\circ})$ when there are $N_{u}$ users in a cell with $120^{\circ}$ horizontal coverage and $30^{\circ}$ vertical coverage. Note that we subject the output of the functions, phase, and delay values to quantization before illustrating their results. We present our findings of both analytical (Sec. IV-A) and iterative (Sec. IV-B) results. Also, we present a run-time comparison between our proposed method and state-of-the-art [5].
IV-A Results of analytical algorithms
@@
In Secs. III-A and III-B, we present our analytical methods to configure PS and TTD units. In Fig. 2(a), we illustrate the average gains of the users, which is denoted as $G_{l}$ in (5), for $N_{u}=2$, where User 1 gets allocated $\alpha$ of the total BW (${\rm BW}$), and User 2 gets $1-\alpha$ of the BW. Our findings reveal that joint optimization consistently outperforms separated optimization by a margin ranging from $0.6$ dB to $4$ dB, depending on the BW allocation ratio. The mean difference between joint and separated optimization, observed across various BW allocations, is approximately $1.29$ dB when the method is least squares (LS), which is an extension of [10] from 2D to 3D, and $1.18$ dB when the method is infinity norm. Hence, it can be deduced that the greater number of systems of equations in the analytical derivation yields higher beamforming gain at the price of higher complexity.
Moreover, the infinity-norm solution results in a higher average gain than LS for unequal BW allocation ($\alpha\leq 0.2$) in both joint and separated optimization scenarios. Although it has a higher computational complexity, the infinity-norm solution treats the two users more fairly and can achieve a larger sum data rate.
@@
In Fig. 2(b), we demonstrate $G_{{\rm mean},i}$ for $i^{\rm th}$ user, defined in (4), when $N_{u}=5$. The BW allocations of the users are $\bm{\alpha}=[0.3,0.2,0.15,0.1,0.25]$. In joint optimization, the infinity norm method yields a maximum gain difference of $3$ dB among users, whereas the LS method results in a $5$ dB difference. A similar trend is observed in separated optimization, with the infinity norm leading to a maximum difference of $3.2$ dB, while LS results in a $5.4$ dB difference. This suggests that the infinity norm method fosters greater fairness in user gain distribution, even when their assigned ratios are unequal. Furthermore, due to the more equitable distribution of average gains among users, the infinity norm also elevates the average gain of users, as illustrated in Fig. 2(a).
@@
Furthermore, we illustrate the impact of joint and separated optimization on beamforming gain in Fig. 3 when the solution is obtained by LS. Here, we demonstrate the maximum gain among subcarriers for every azimuth and elevation angle when the desired user locations are marked with red diamonds. More sidelobes are observed in the case of separated optimization, as depicted in Fig. 3(a), in contrast to the joint optimization shown in Fig. 3(b). This difference can be attributed to the interdependence of the phase and delay elements, which is fully captured in the joint analytical solution. In Fig. 4, we display the gain across various subcarriers while maintaining a constant elevation angle, which corresponds to the User 3 in Fig. 3(b) with $(\theta_{\rm az},\theta_{\rm el})=(0^{\circ},105^{\circ})$. Here, the red line is the desired maximum gain allocation from -9.5 MHz to 9.5 MHz, which is fairly satisfied by using joint optimization.
IV-B Results of iterative algorithms